# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ 3.1 Requirements motivated by the German DiGA guide
As an illustrative example, we consider a case inspired by the DiGA elevida, which targets fatigue in patients with multiple sclerosis and has been permanently registered in the German DiGA directory via the fast-track procedure (see [16]). The study [17] for permanent registration was a parallel group, two-arm RCT. The primary outcome was the physical and cognitive fatigue measured on the Chalder Fatigue Scale. The healthcare effect was defined by the mean difference between a waitlist control group and the treatment group (using DiGA for 12 weeks). As mentioned in [17], the minimally important difference for improvement on the Chalder Fatigue Scale was estimated between $0.7$ and $1.4$. In the following, we will work with $\delta_{\text{rel}}=1$. We assume hypothetically that a fast-track procedure is planned for a similar DiGA by another applicant, based on the information of the DiGA elevida. In this fast-track procedure, we use parallel group designs for both, conditional and permanent registration, each with balanced samples and normal mean differences as estimates for the healthcare effect. The informations for the first and second stage are then given by $I_{k}=n_{k}/(2\sigma^{2})$ where $n_{k}\in\mathbb{N}$ is the sample size per group for stage $k$, $k=1,2$. A sample size calculation with $\alpha=0.025$ and $0.81-\beta=$ 0.8 at $\delta=\delta_{\text{rel}}=1$ would give the information $I_{\text{rel}}=2.8^{2}/1^{2}=7.84$. Using the estimated variance $\hat{\sigma}^{2}=5.17^{2}$ from [17], the corresponding per-group sample size would be $n_{1,\text{rel}}:=420\approx 2\hat{\sigma}^{2}I_{\text{rel}}$. This implies for example that a pilot study with $t_{\text{rel}}(I_{1})=0.5$ corresponds to a study with per-group sample size of $n_{1}=0.5\cdot 420=210$.
@@
$I_{1}<I_{1,\max}:={\Phi^{-1}(1-\alpha)^{2}}/{\delta_{\text{rel}}^{2}}\leavevmode\nobreak\ .$ (2)
A pilot study with information $I_{1}\geq I_{1,\max}$ and the requirement $p_{1}\leq\alpha_{f}\ (\leq\alpha)$ wouldn’t make sense for the applicants, because it would be less or as efficient as a similarly large study for permanent registration for which only $p_{1}\leq\alpha$ is required. Therefore we only consider pilot information $I_{1}$ that are lower than the universal upper bound $I_{1,\max}$.
@@ 3.2 Consequences in conjunction with required conditional registration
This subsection analyses the consequences of the requirements for a conditional registration for situations where a unsuccessful conditional registration renders a study for a permanent registration impossible (for whatever reasons), i.e. where conditional registration is required. A justified aim of the applicants, but also regulators, should be to guarantee a certain probability of overall success, which we understand as the probability of achieving permanent registration. This leads to the requirement that the success probability of the conditional registration needs to be sufficiently large. For example, to reach an overall success probability of at least $1-\beta$ at some $\delta\geq\delta_{\text{rel}}$, we must have a pilot study with "power"
$\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f}\right)>1-\beta\leavevmode\nobreak\ ,$ (3)
@@
Note at first, that this condition cannot be fulfilled (for $0.51-\beta>$ 0.5) if $\delta=\delta_{\text{rel}}$: In this case, the power in (3) cannot exceed $0.5$ because $Z_{1}\geq z_{f}$ implies $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$. So if $\delta=\delta_{\text{rel}}$ is considered highly likely and worthwhile for registration, conducting a single trial for permanent registration is more promising than a fast-track registration process with required conditional registration.
@@
$\xi:={\delta}/{\delta_{\text{rel}}}\in(1,\infty)\leavevmode\nobreak\ .$
Figure 2(a) shows, in dependence on $\xi$, the universal lower bound (5) in the relative scale $t_{\text{rel}}(I_{1,\min})={I_{1,\min}}/{I_{\text{rel}}}$. Note that $t_{\text{rel}}(I_{1,\min})$ is independent from $\delta_{\text{rel}}$ and only depends on $\alpha$, $\beta$, $\alpha_{c}$ and $\xi$. The first part of the maximum in (5) comes from the requirement $p_{1}\leq\alpha_{\text{rel}}$ and is represented in Figure 2(a) by the black solid decreasing line. The other lines represent the second part of the maximum in (5) (which comes from the requirement $p_{1}\leq\alpha_{c}$) for the different (one-sided) levels $\alpha_{c}=0.05,0.1,0.15$. The minimum possible relative Information $t_{\text{rel}}(I_{1,\min})$ can also be read off Figure 2(a) by using for each $\alpha_{c}$ the maximum of the corresponding horizontal line and the solid black line.
The universal upper bound $I_{1,\max}$ for the information of a reasonable pilot study is also shown in Figure 2(a) on the relative scale, i.e. $t_{\text{rel}}(I_{1,\max})$, namely by the horizontal grey dot-dashed line. Recall that this universal upper comes from the regulatory requirement $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ and the reflection that the requirements for conditional registration should not be more stringent than for a permanent registration. One can see that the universal lower bound $t_{\text{rel}}(I_{1,\min})$ can exceed this universal upper bound in which case the fast-track procedure with a required conditional registration does not make sense. This is the case for $\xi$ falling short of the intersection point between the decreasing black solid line and grey dot-dashed line, which equals
@@
With $\alpha=0.025$ and $\beta=0.2$ we obtain $\xi_{\min}\approx 1.43$. This implies that the here considered fast-track process is useful only, if we can assume a health care effect that is at least 43% higher than the minimal relevant effect $\delta_{\text{rel}}$. Note that this effect would even be larger with a larger conditional power of e.g. $0.91-\beta=$ 0.9. All the following considerations of the fast-track procedure with required conditional registration will be made under the assumptions $\alpha=0.025$ and $\beta=0.2$.
(a) Universal lower bound for $t_{\text{rel}}(I_{1})={I_{1}}/{I_{\text{rel}}}$ in
dependence on $\xi=\delta/\delta_{\text{rel}}\geq\xi_{\min}=1.43$.
(b) Universal lower bound for $t_{\xi}(I_{1})={I_{1}}/{I_{\delta}}$ in
dependence on $\xi=\delta/\delta_{\text{rel}}\geq\xi_{\min}=1.43$.
Figure 2: Visualization of the minimum needed information for the first period of data collection.
@@
Since the observed healthcare effect in [17] was $2.74$, a reasonable and conservative a priori healthcare effect assumption is $\delta=2=2\cdot\delta_{\text{rel}}$, i.e. $\xi=2$. We can read off Figure 2(a), that $t_{\text{rel}}(I_{1,\min})=I_{1,\min}/I_{\text{rel}}=n_{1,\min}/n_{1,\text{rel}}=0.2$ for the choice of $\alpha_{c}=0.05$. This implies that, to meet (3), the fast-track procedure needs a minimum first stage sample size per group of $n_{1,\min}=84=0.2\cdot 420$. Under the assumptions a fixed-size sample test with significance level $\alpha=0.025$ and power $0.81-\beta=$ 0.8 would need the information of $I_{\delta}:=\eta_{f}^{2}/\delta^{2}=2.8^{2}/2^{2}=1.96$, i.e. a per group sample size of $n_{\delta}:=105\approx 2\hat{\sigma}^{2}I_{\delta}$.
Next, we discuss the choice of $\alpha_{c}$ for the fast-track procedure with required conditional registration with reference to Figure 2(b). This figure shows the universal lower bound for the pilot information on another scale, namely relative to the information $I_{\delta}:={\eta_{f}^{2}}/{\delta^{2}}$ needed for achieving a power of $1-\beta$ at the true or a priori assumed alternative $\vartheta=\delta$ with a single fixed-size sample test: The vertical axis of this figure is
@@
and it shows $t_{\xi}(I_{1,\min})=t_{\text{rel}}(I_{1,\min})\xi^{2}$ in dependence of $\xi$. Since this scale increases with $\xi^{2}$, the horizontal grey line for the universal upper bound in Figure 2(a) becomes a curved line in Figure 2(b), and the non-solid curved lines in Figure 2(a), representing the part of the universal lower bound coming from the requirement $p_{1}\leq\alpha_{c}$, are becoming straight lines in Figure 2(a). The universal lower bound for the pilot information (to achieve a power larger than $1-\beta$) is again represented by the maximum of the solid black curve and horizontal line corresponding to the respective $\alpha_{c}$. One can see from this figure that with $\alpha_{c}=0.05$ the minimum information for the pilot study must exceed $78\%\ (\approx 80\%)$ of the information $I_{\delta}$ required for a fixed size sample test with a power of $1-\beta$ at the true or a priori assumed alternative $\vartheta=\delta$. This indicates that $\alpha_{c}=0.05$ may lead to rather large pilot studies and that a lager choice for $\alpha_{c}$ would be more practical. Since the level $\alpha_{c}=0.15$ leads to a universal lower bound below $I_{\delta}/2$ (more precisely, $44,94\%$ of $I_{\delta}$) and may still be considered an acceptable level for conditional registration, we will use this value in our further investigations. Recall that for our example fast-track procedure we have for $\delta=2$ that $I_{\delta}=1.96$ with corresponding per group sample size $n_{\delta}=105$. From Figure 2(b) for $\alpha_{c}=0.15$ it follows that $0.45=t_{\xi}(I_{1,\min})=n_{1,\min}/n_{\delta}$. That means $n_{1,\min}=48\approx 0.45\cdot n_{\delta}$ is now the minimum needed first-stage sample size per group and significantly smaller than $84$, which is the one needed if $\alpha_{c}$ is not $0.15$ but $0.05$.
Let us further explain, why we considered the two different scales $t_{\text{rel}}$ and $t_{\xi}$: Depending on the applicant’s stance, i.e. the confidence in the assumed a priori effect $\delta$, one scale may be more relevant than the other. If there is little confidence in the assumed a priori effect $\delta$ and few or no data has been collected before starting collecting data as part of the fast-track procedure, the $t_{\text{rel}}$-scale may be more interesting. However, if applicants have greater confidence in the correctness of the assumed a priori effect $\delta$, comparing the information with the one needed for a fixed-size sample test powered at $\delta$ may be more appropriate. In this case, applicants especially do not likely want the (minimum) information for the first stage to exceed $I_{\delta}$, i.e. $t_{\xi}(I_{1,\min})<1$ should hold.
@@ 4 Study designs with required conditional registration
In this section, we assume that at the time of the planning of a fast-track procedure by the applicants, an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration. Reasons for this may be regulatory requirements or financial aspects, for example. We relax this assumption in Section 5. We will investigate the efficiency of the standard design with two separate studies when a certain overall power is anticipated, and compare them with adaptive design after a brief recap of adaptive designs.
@@ 4.1 Overall performance of the standard fast-track registration process
We consider the overall and second stage power and corresponding sample sizes of the standard design with two separate stages (studies). The overall power calculation (including both stages) is done for some a priori assumed effect $\delta$ with $\xi=\delta/\delta_{\text{rel}}>\xi_{\min}$ for the power $1-\beta$. We further assume that, after a successful first stage, the second stage is powered for the same value $1-\beta$, but now at the observed first stage estimate $\hat{\vartheta}_{1}$, using the well-known sample size formula for z-tests. Recall that a successful first stage implies $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$.
Since the calculated second stage information can become arbitrarily small for large interim effect estimates, but the second stage study should to be sufficiently large to be sufficiently convincing for an permanent registration, we also assume a specific minimum information $I_{2,\min}$ that is never undercut, even if the second stage sample size calculation falls short of this minimum. The selection of this minimum stage-two information $I_{2,\min}$ also provides a way to control the overall power with a fixed stage-one information. Fixing the stage-one information and choosing $I_{2,\min}$ to reach a sufficiently large overall success probability may be of particular interest for small (start-up) companies that have a relatively small budget for the pilot study but can recruit additional funds for the second study during the temporary registration period. In the following, we will consider the minimal information $I_{2,\min}$ required for an overall success probability ($1-\beta$) as a function of a (given) first-stage information $I_{1}$ assuming a specific health care effect $\delta=\xi\delta_{\text{rel}}$ for $\xi>\xi_{\min}$. This will inform us about the efficiency of the standard fast track procedure under different assumptions and to compare it with alternative designs utilizing adaptive design methodology.
As described, we assume that the fast-track procedure leads to a successful permanent registration if the first stage is successful (with conditional registration) and, at the second stage, the hypothesis $H_{0}$ is rejected by the z-test with second stage p-value $p_{2}\leq\alpha$. Therefore, the overall power can be written as
[…]
@@
$n_{2}(z_{1}):=2\hat{\sigma}^{2}I_{2}(z_{1})=\max\left\{n_{2,\min},2\hat{\sigma}^{2}\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\}^{2}/\hat{\vartheta}_{1}^{2}\right\}.$
@@
Recall from Section 3, that to achieve the overall power $1-\beta$ at the assumed a priori effect $\delta=\xi\delta_{\text{rel}}$ for $\xi>\xi_{\min}$, the information of the first stage must be chosen greater than $I_{1,\min}$ defined in (5). This results from the fact that the overall power is less than the power for a successful conditional registration. It can be shown that for $\xi>\xi_{\min}$ there exists for all $I_{1}>I_{1,\min}$ a unique minimum information $I_{2,\min}$ such that $\mathcal{S}(I_{2,\min},I_{1})=1-\beta$ if $\mathcal{S}(0,I_{1})\leq 1-\beta$, and if $\mathcal{S}(0,I_{1})>1-\beta$ the overall power is achieved without a minimum information for the second stage, in which case we set $I_{2,\min}=0$. The proof of this result can be be found in Appendix A.
Again, for a more universal investigation, we consider relative informations rather than the absolute informations $I_{2,\min}$ and $I_{1}$. Since we belief, that in practice, the fixed-size sample test for the a priori assumed $\delta=\xi\delta_{\text{rel}}$ is the more relevant benchmark than the fixed-size sample test for $\delta_{\text{rel}}$, we will present the results on the $t_{\xi}$-scale, however, now for a fixed $\xi$. Because
@@
The red dotted lines in Figure 3 show for $\xi=1.75$ and $\xi=2$ in dependence on $t_{\xi}(I_{1})$, where $t_{\xi}(I_{1,\min})<t_{\xi}(I_{1})\leq t_{\xi}(I_{1,\max})$, the minimum relative information $t_{\xi}(I_{2,\min})$ required with the standard fast track process with required conditional registration to have overall power $0.81-\beta=$ 0.8. The other lines will be described and discussed in the next section. The R-code used for the calculation of $t_{\xi}(I_{2,\min})$ and for all subsequent images can be downloaded from https://github.com/LianeKluge/Adaptive-Designs-in-Fast-Track-Registration-Processes. As it can be seen from Figure 3, when $I_{1}$ comes close to the critical value $I_{1,\min}$, the second stage information becomes large. This is because when $I_{1}$ comes close to the critical value, the power for the conditional registration comes close to $1-\beta$. Then, the second stage information required for the anticipated overall power becomes larger to compensate for the requirement to also reject $H_{0}$ at the second stage at level $\alpha$. This behaviour applies to any design for which $Z_{1}\geq z_{f}$ is a requirement for a final success. We note that for the standard design with two separate stages, the requirement for a specific overall power demands that the minimum information for the second stage is always greater than $I_{\delta}$, i.e. the sample size required for a fixed-size sample test with a power $1-\beta$ at the a priori assumed effect $\delta$. For our example fast-track procedure we have for $\xi=2$ due to the choice of $n_{1}=63$ that $t_{\xi}(I_{1})=n_{1}/n_{\delta}=63/105=0.6$. According to Figure 3(b) the minimum needed relative information for the second stage corresponding to $t_{\xi}(I_{1})=0.6$ is $1.41=t_{\xi}(I_{2,\min})=n_{2,\min}/n_{\delta}$. That means the minimum needed per group sample size for the second stage is $n_{2,\min}=1.41\cdot 105=148$.
In the next subsection, we will discuss adaptive designs for fast-track procedures, which have some important advantages. Compared to the non-adaptive design, where the overall level of achieving both the conditional and the permanent registration is $\alpha_{f}\cdot\alpha$, the overall level is $\alpha$ for adaptive designs. This leads to greater efficiency when aiming to achieve a certain overall power. Moreover, in contrast to the standard design with separate studies, the data from the first stage can also be used for the final confirmatory hypothesis test of $H_{0}$. This increases efficiency even more. The performance of two adaptive designs is teased in Figure 3 by the black dashed line and the blue solid line.
(a) $\xi=1.75$.
(b) $\xi=2$.
Figure 3: Minimum needed information for the second stage to achieve overall power of $0.81-\beta=$ 0.8 for the common non-adaptive design and the adaptive designs, inverse normal method and Fisher’s product test.

### v3

@@ 3.1 Requirements motivated by the German DiGA guide
As an illustrative example, we consider a case inspired by the DiGA elevida, which targets fatigue in patients with multiple sclerosis and has been permanently registered in the German DiGA directory via the fast-track procedure [22]. The study for permanent registration was a parallel group, two-arm RCT [23]. The primary outcome was the physical and cognitive fatigue measured on the Chalder Fatigue Scale. The healthcare effect was defined by the mean difference between a waitlist control group and the treatment group (using DiGA for 12 weeks). As mentioned in [23], the minimally important difference for improvement on the Chalder Fatigue Scale was estimated between $0.7$ and $1.4$. In the following, we will work with $\delta_{\text{rel}}=1$. We assume hypothetically that a fast-track procedure is planned for a similar DiGA by another applicant, based on the information of the DiGA elevida. In this fast-track procedure, we use parallel group designs for both conditional and permanent registration, each with balanced samples and normal mean differences as estimates for the healthcare effect. The informations for the first and second stage are then given by $I_{k}=n_{k}/(2\sigma^{2})$ where $n_{k}\in\mathbb{N}$ is the sample size per group for stage $k$, $k=1,2$. A sample size calculation with $\alpha=0.025$ and $1-\beta=0.8$ at $\delta_{\text{rel}}=1$ would give the information $I_{\text{rel}}=2.8^{2}/1^{2}=7.84$. Using the estimated variance $\hat{\sigma}^{2}=5.17^{2}$ from [23], the corresponding per-group sample size would be $n_{1,\text{rel}}:=420\approx 2\hat{\sigma}^{2}I_{\text{rel}}$. This implies, for example, that a pilot study with $t_{\text{rel}}(I_{1})=0.5$ corresponds to a study with per-group sample size of $n_{1}=0.5\cdot 420=210$.
@@
$I_{1}<I_{1,\max}:={\Phi^{-1}(1-\alpha)^{2}}/{\delta_{\text{rel}}^{2}}\penalty 10000\ .$ (2)
A pilot study with information $I_{1}\geq I_{1,\max}$ and the requirement $p_{1}\leq\alpha_{f}\ (\leq\alpha)$ wouldn’t make sense for the applicants, because it would be less or as efficient as a similarly large study for permanent registration for which only $p_{1}\leq\alpha$ is required. Therefore, we only consider pilot information $I_{1}$ that is lower than the universal upper bound $I_{1,\max}$.
@@ 3.2 Consequences in conjunction with required conditional registration
This subsection analyses the consequences of the requirements for a conditional registration for situations where an unsuccessful conditional registration renders a study for a permanent registration impossible (for whatever reasons), i.e. where conditional registration is required. A justified aim of the applicants, but also regulators, should be to guarantee a certain probability of overall success, which we understand as the probability of achieving permanent registration. This leads to the requirement that the success probability of the conditional registration needs to be sufficiently large. For example, to reach an overall success probability of at least $1-\beta$ at some $\delta\geq\delta_{\text{rel}}$, we must have a pilot study with “power”
$\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f}\right)>1-\beta\penalty 10000\ ,$ (3)
@@
Note at first that this condition cannot be fulfilled (for $1-\beta>0.5$) if $\delta=\delta_{\text{rel}}$: In this case, the power in (3) cannot exceed $0.5$ because $Z_{1}\geq z_{f}$ implies $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$. So if $\delta=\delta_{\text{rel}}$ is considered highly likely and worthwhile for registration, conducting a single trial for permanent registration is more promising than a fast-track registration process with required conditional registration.
@@
$\xi:={\delta}/{\delta_{\text{rel}}}\in(1,\infty)\penalty 10000\ .$
Figure 2 (left) shows, in dependence on $\xi$, the universal lower bound (5) in the relative scale $t_{\text{rel}}(I_{1,\min})={I_{1,\min}}/{I_{\text{rel}}}$. Note that $t_{\text{rel}}(I_{1,\min})$ is independent from $\delta_{\text{rel}}$ and only depends on $\alpha$, $\beta$, $\alpha_{c}$ and $\xi$. The first part of the maximum in (5) comes from the requirement $p_{1}\leq\alpha_{\text{rel}}$ and is represented in Figure 2 (left) by the solid decreasing line. The other lines represent the second part of the maximum in (5) (which comes from the requirement $p_{1}\leq\alpha_{c}$) for the different (one-sided) levels $\alpha_{c}=0.05,0.1,0.15$. The minimum possible relative Information $t_{\text{rel}}(I_{1,\min})$ can also be read off Figure 2 (left) by using for each $\alpha_{c}$ the maximum of the corresponding horizontal line and the decreasing solid line.
The universal upper bound $I_{1,\max}$ for the information of a reasonable pilot study is also shown in Figure 2 (left) on the relative scale, i.e. $t_{\text{rel}}(I_{1,\max})$, namely by the horizontal dot-dashed line. Recall that this universal upper comes from the regulatory requirement $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ and the reflection that the requirements for conditional registration should not be more stringent than for a permanent registration. One can see that the universal lower bound $t_{\text{rel}}(I_{1,\min})$ can exceed this universal upper bound, in which case the fast-track procedure with a required conditional registration does not make sense. This is the case for $\xi$ falling short of the intersection point between the decreasing solid line and dot-dashed line, which equals
@@
With $\alpha=0.025$ and $\beta=0.2$ we obtain $\xi_{\min}\approx 1.43$. This implies that the here considered fast-track process is useful only if we can assume a health care effect that is at least 43% higher than the minimal relevant effect $\delta_{\text{rel}}$. Note that this effect would even be larger with a larger overall power of e.g. $1-\beta=0.9$. All the following considerations of the fast-track procedure with required conditional registration will be made under the assumptions $\alpha=0.025$ and $\beta=0.2$.
Figure 2: Visualisation of the minimum information needed for the first period of data collection. The universal lower bound for $t_{\text{rel}}(I_{1})={I_{1}}/{I_{\text{rel}}}$ in dependence on $\xi=\delta/\delta_{\text{rel}}\geq\xi_{\min}=1.43$ (left) and the universal lower bound for $t_{\xi}(I_{1})={I_{1}}/{I_{\delta}}$ in dependence on $\xi\geq\xi_{\min}$ (right) are shown.
@@
Since the observed healthcare effect in the conducted trial [23] was $2.74$, a reasonable and conservative a priori healthcare effect assumption is $\delta=2=2\cdot\delta_{\text{rel}}$, i.e. $\xi=2$. We can read off Figure 2 (left), that $t_{\text{rel}}(I_{1,\min})=I_{1,\min}/I_{\text{rel}}=n_{1,\min}/n_{1,\text{rel}}=0.2$ for the choice of $\alpha_{c}=0.05$. This implies that, to meet (3), the fast-track procedure needs a minimum first stage sample size per group of $n_{1,\min}=84=0.2\cdot 420$. For comparison, a fixed-size sample test with significance level $\alpha=0.025$ and power $1-\beta=0.8$ at $\delta$ would need the information of $I_{\delta}:=\eta_{f}^{2}/\delta^{2}=2.8^{2}/2^{2}=1.96$, i.e. a per group sample size of $n_{\delta}:=105\approx 2\hat{\sigma}^{2}I_{\delta}$.
Next, we discuss the choice of $\alpha_{c}$ for the fast-track procedure with required conditional registration with reference to Figure 2 (right). This figure shows the universal lower bound for the pilot information on another scale, namely relative to the information $I_{\delta}:={\eta_{f}^{2}}/{\delta^{2}}$ needed for achieving a power of $1-\beta$ at the true or a priori assumed alternative $\vartheta=\delta$ with a single fixed-size sample test: The vertical axis of this figure is
@@
and it shows $t_{\xi}(I_{1,\min})=t_{\text{rel}}(I_{1,\min})\xi^{2}$ in dependence of $\xi$. Since this scale increases with $\xi^{2}$, the horizontal dotted-dashed line for the universal upper bound in Figure 2 (left) becomes a curved line in Figure 2 (right), and the non-solid curved lines in Figure 2 (left), representing the part of the universal lower bound coming from the requirement $p_{1}\leq\alpha_{c}$, are becoming straight lines in Figure 2 (left). The universal lower bound for the pilot information (to achieve a power larger than $1-\beta$) is again represented by the maximum of the decreasing solid curve and the horizontal line corresponding to the respective $\alpha_{c}$. One can see from this figure that with $\alpha_{c}=0.05$ the minimum information for the pilot study must exceed $78\%\ (\approx 80\%)$ of the information $I_{\delta}$ required for a fixed-size sample test with a power of $1-\beta$ at the true or a priori assumed alternative $\vartheta=\delta$. This indicates that $\alpha_{c}=0.05$ may lead to rather large pilot studies and that a larger choice for $\alpha_{c}$ would be more practical. Since the level $\alpha_{c}=0.15$ leads to a universal lower bound below $I_{\delta}/2$ (more precisely, $44,94\%$ of $I_{\delta}$) and may still be considered an acceptable level for conditional registration, we will use this value in our further investigations. Recall that for our example fast-track procedure, we have for $\delta=2$ that $I_{\delta}=1.96$ with corresponding per group sample size $n_{\delta}=105$. From Figure 2 (right) for $\alpha_{c}=0.15$ it follows that $0.45=t_{\xi}(I_{1,\min})=n_{1,\min}/n_{\delta}$. That means $n_{1,\min}=48\approx 0.45\cdot n_{\delta}$ is now the minimum needed first-stage sample size per group and significantly smaller than $84$, which is the one needed if $\alpha_{c}$ is not $0.15$ but $0.05$.
Let us further explain why we considered the two different scales $t_{\text{rel}}$ and $t_{\xi}$: Depending on the applicant’s stance, i.e. the confidence in the assumed a priori effect $\delta$, one scale may be more relevant than the other. If there is little confidence in the assumed a priori effect $\delta$ and few or no data have been collected before starting to collect data as part of the fast-track procedure, the $t_{\text{rel}}$-scale may be more interesting. However, if applicants have greater confidence in the correctness of the assumed a priori effect $\delta$, comparing the information with the one needed for a fixed-size sample test powered at $\delta$ may be more appropriate. In this case, applicants especially do not likely want the (minimum) information for the first stage to exceed $I_{\delta}$, i.e. $t_{\xi}(I_{1,\min})<1$ should hold.
@@ 4 Study designs with required conditional registration
In this section, we assume that at the time of the planning of a fast-track procedure by the applicants, an unsuccessful study for conditional registration precludes any further attempt to receive a permanent registration. Reasons for this may be regulatory requirements or financial aspects, for example. We relax this assumption in Section 5. We will investigate the efficiency of the standard design with two separate studies when a certain overall power is anticipated, and compare them with adaptive designs after a brief recap of adaptive designs.
@@ 4.1 Overall performance of the standard fast-track registration process
We consider the overall and second-stage power and corresponding sample sizes of the standard design with two separate stages (studies). The overall power calculation (including both stages) is done for some a priori assumed effect $\delta$ with $\xi=\delta/\delta_{\text{rel}}>\xi_{\min}$ for the power $1-\beta$. We further assume that, after a successful first stage, the second stage is powered for the same value $1-\beta$, but now at the observed first stage estimate $\hat{\vartheta}_{1}$, using the well-known sample size formula for z-tests. Recall that a successful first stage implies $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$.
Since the calculated second stage information can become arbitrarily small for large interim effect estimates, but the second stage study should to be sufficiently large to be sufficiently convincing for a permanent registration, we also assume a specific minimum information $I_{2,\min}$ that is never undercut, even if the second stage sample size calculation falls short of this minimum. The selection of this minimum stage-two information $I_{2,\min}$ also provides a way to control the overall power with a fixed stage-one information. In the following, we will consider the minimal information $I_{2,\min}$ required for an overall success probability ($1-\beta$) as a function of a (given) first-stage information $I_{1}$ assuming a specific health care effect $\delta=\xi\delta_{\text{rel}}$ for $\xi>\xi_{\min}$. This will inform us about the efficiency of the standard fast-track procedure under different assumptions and compare it with alternative designs utilizing adaptive design methodology.
As described, we assume that the fast-track procedure leads to a successful permanent registration if the first stage is successful (with conditional registration) and, at the second stage, the hypothesis $H_{0}$ is rejected by the z-test with second-stage p-value $p_{2}\leq\alpha$. Therefore, the overall power can be written as
$\mathcal{S}(I_{2,\min},I_{1}):=\,$ $\mathbb{P}_{\delta}\left(Z_{1}\geq z_{f},p_{2}\leq\alpha\right)=\int\limits_{z_{f}}^{\infty}\mathbb{P}_{\delta}(p_{2}\leq\alpha\penalty 10000\ |\penalty 10000\ Z_{1}=z_{1})\,\phi\big(z_{1}-\sqrt{I}_{1}\delta\big)dz_{1}$
$=\int\limits_{z_{f}}^{\infty}\left\{1-\Phi\left(\Phi^{-1}(1-\alpha)-\sqrt{\max\big\{I_{2,\min},I_{1}\eta_{f}^{2}/z_{1}^{2}\big\}}\delta\right)\right\}\phi\big(z_{1}-\sqrt{I}_{1}\delta\big)dz_{1},$ (7)
@@
$n_{2}(z_{1}):=2\hat{\sigma}^{2}I_{2}(z_{1})=\max\left\{n_{2,\min},2\hat{\sigma}^{2}\left\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\right\}^{2}/\hat{\vartheta}_{1}^{2}\right\}.$ (9)
@@
Recall from Section 3 that to achieve the overall power $1-\beta$ at the assumed a priori effect $\delta=\xi\delta_{\text{rel}}$ for $\xi>\xi_{\min}$, the information of the first stage must be chosen greater than $I_{1,\min}$ defined in (5). This results from the fact that the overall power is less than the power for a successful conditional registration. It can be shown that for $\xi>\xi_{\min}$ there exists for all $I_{1}>I_{1,\min}$ a unique minimum information $I_{2,\min}$ such that $\mathcal{S}(I_{2,\min},I_{1})=1-\beta$ if $\mathcal{S}(0,I_{1})\leq 1-\beta$, and if $\mathcal{S}(0,I_{1})>1-\beta$ the overall power is achieved without a minimum information for the second stage, in which case we set $I_{2,\min}=0$. The proof of this result can be found in the Appendix.
Again, for a more universal investigation, we consider relative informations rather than the absolute informations $I_{2,\min}$ and $I_{1}$. Since we believe that in practice, the fixed-size sample test for the a priori assumed $\delta=\xi\delta_{\text{rel}}$ is the more relevant benchmark than the fixed-size sample test for $\delta_{\text{rel}}$, we will present the results on the $t_{\xi}$-scale, however, now for a fixed $\xi$. Because
@@
The dotted lines in Figure 3 show for $\xi=1.75$ and $\xi=2$ in dependence on $t_{\xi}(I_{1})$, where $t_{\xi}(I_{1,\min})<t_{\xi}(I_{1})\leq t_{\xi}(I_{1,\max})$, the minimum relative information $t_{\xi}(I_{2,\min})$ required with the standard fast track process with required conditional registration to have overall power $1-\beta=0.8$. The other lines will be described and discussed in the next section. The R-code used for the calculation of $t_{\xi}(I_{2,\min})$ and for all subsequent images can be downloaded from https://github.com/LianeKluge/Adaptive-Designs-in-Fast-Track-Registration-Processes. As it can be seen from Figure 3, when $I_{1}$ comes close to the critical value $I_{1,\min}$, the second stage information becomes large. This is because when $I_{1}$ comes close to the critical value, the power for the conditional registration comes close to $1-\beta$. Then, the second stage information required for the anticipated overall power becomes larger to compensate for the requirement to also reject $H_{0}$ at the second stage at level $\alpha$. This behavior applies to any design for which $Z_{1}\geq z_{f}$ is a requirement for a final success. We note that for the standard design with two separate stages, the requirement for a specific overall power demands that the minimum information for the second stage is always greater than $I_{\delta}$, i.e. the sample size required for a fixed-size sample test with a power $1-\beta$ at the a priori assumed effect $\delta$. For our example fast-track procedure we have for $\xi=2$ due to the choice of $n_{1}=63$ that $t_{\xi}(I_{1})=n_{1}/n_{\delta}=63/105=0.6$. According to Figure 3 (right) the minimum needed relative information for the second stage corresponding to $t_{\xi}(I_{1})=0.6$ is $1.41=t_{\xi}(I_{2,\min})=n_{2,\min}/n_{\delta}$. That means the minimum needed per group sample size for the second stage is $n_{2,\min}=1.41\cdot 105=148$.
In the next subsection, we will discuss adaptive designs for fast-track procedures, which have some important advantages. Compared to the non-adaptive design, where the overall level of achieving both the conditional and the permanent registration is $\alpha_{f}\cdot\alpha$, the overall level is $\alpha$ for adaptive designs. This leads to greater efficiency when aiming to achieve a certain overall power. Moreover, in contrast to the standard design with separate studies, the data from the first stage can also be used for the final confirmatory hypothesis test of $H_{0}$. This increases efficiency even more. The performance of two adaptive designs is teased in Figure 3 by the dashed line and the solid line.
Figure 3: Minimum needed information for the second stage to achieve overall power of $1-\beta=0.8$ for $\xi=1.75$ (left) and $\xi=2$ (right) for the common non-adaptive design and the adaptive designs, inverse normal method and Fisher’s product test.